# CaLiV: LiDAR-to-Vehicle Calibration of Arbitrary Sensor Setups via Object Reconstruction

### v1

@@ Abstract
In autonomous systems, sensor calibration is essential for a safe and efficient navigation in dynamic environments. Accurate calibration is a prerequisite for reliable perception and planning tasks such as object detection and obstacle avoidance. Many existing LiDAR calibration methods require overlapping fields of view, while others use external sensing devices or postulate a feature-rich environment. In addition, Sensor-to-Vehicle calibration is not supported by the vast majority of calibration algorithms. In this work, we propose a novel target-based technique for extrinsic Sensor-to-Sensor and Sensor-to-Vehicle calibration of multi-LiDAR systems called CaLiV. This algorithm works for non-overlapping FoVs, as well as arbitrary calibration targets, and does not require any external sensing devices. First, we apply motion to produce FoV overlaps and utilize a simple unscented Kalman filter to obtain vehicle poses. Then, we use the Gaussian mixture model-based registration framework GMMCalib to align the point clouds in a common calibration frame. Finally, we reduce the task of recovering the sensor extrinsics to a minimization problem. We show that both translational and rotational Sensor-to-Sensor errors can be solved accurately by our method. In addition, all Sensor-to-Vehicle rotation angles can also be calibrated with high accuracy. We validate the simulation results in real-world experiments. The code is open source and available on https://github.com/TUMFTM/CaLiV.
@@ I INTRODUCTION
A similar problem occurs in sensor calibration when dealing with non-overlapping sensor setups. The calibration becomes significantly more challenging with sensors that do not share a common field of view and, therefore, never simultaneously observe the same features. This is common in applications involving large vehicles, such as buses or trains. Accordingly, non-overlapping sensor calibration frameworks also use motion-based methods, including SLAM or hand-eye calibration.
SLAM-based calibration methods generate a map of the environment through point cloud matching and implicitly solve for the extrinsic parameters. These approaches rely heavily on environmental features, which may not always be available. Hand-eye calibration [2] methods are less dependent on specific environmental features, but do not achieve the calibration accuracy of SLAM algorithms [3]. In addition, these motion-based methods are commonly presented as targetless calibration techniques [4, 5]. Although these targetless approaches offer greater flexibility, they still do not match the accuracy of target-based solutions [6]. In contrast, target-based S2V or non-overlapping S2S calibration has mainly been addressed using external sensing devices [7, 8, 9].
Figure 1: CaLiV: Data is collected by driving a curved trajectory. The data is first aligned using vehicle poses and different perspectives of the calibration target. In the optimization phase, the algorithm first computes the S2S calibration, which serves as input for the S2V optimization. The input to the S2S and S2V optimization is a calibration matrix and the shape of the reconstructed target, both generated using GMMCalib [10], as well as the vehicle poses from an unscented Kalman Filter (UKF).
In this work, we present CaLiV, a two-stage calibration method as a target-based approach to address the problems of S2V calibration for an arbitrary sensor setup without any additional external sensing devices. The first stage aims to perform (non-overlapping) S2S calibration. The second stage utilizes the S2S calibration to compute the S2V calibration parameters. As the underlying alignment model, we rely on GMMCalib [10], which provides a specific reference frame formulation for simultaneous optimization of the S2V calibration parameters. The approach is illustrated in Fig. 1. Our contributions are as follows:
  • To the best of our knowledge, we are the first to present an open-source target-based Sensor-to-Vehicle calibration framework for a multi-LiDAR sensor setup without any external sensor devices.
  • In addition, our approach is the first to simultaneously allow for Sensor-to-Sensor calibration of arbitrary sensor setups.
  • We utilize a joint registration method and incorporate target reconstruction to perform Sensor-to-Vehicle calibration using a reference frame formulation as presented in GMMCalib.
@@ II-B Target-Based Calibration
The state-of-the-art shows that for the above requirements, target-based approaches outperform targetless approaches in terms of calibration accuracy. However, this is mainly due to the use of external sensing devices and very specific calibration targets, which in some cases, as shown in [8], are even measured in advance. Accuracy is therefore achieved at the expense of the flexibility that can be provided by targetless approaches. We conclude that there is no target-based, environment-independent calibration framework for extrinsic S2V calibration of multi-LiDAR systems with non-overlapping FoVs. Therefore, to combine the advantages of both strategies, we propose a target-based calibration framework that makes use of feature-based and motion-based principles for S2V calibration as well as the ability to calibrate non-overlapping sensors. We achieve this with arbitrary targets and without the use of additional external sensor systems.
@@ III METHODOLOGY
This section describes the methodology behind the framework of Fig. 1. For the sake of clarity, we consider the calibration with a vehicle with two non-overlapping LiDARs. However, the proposed technique can easily be generalized for any multi-LiDAR setup. In the following, the notation of [26] is used to express all mathematical formulations related to homogeneous transformations.
III-A Problem Statement
@@
Figure 2: Spatial relationship between the frames $\{V^{i}\}$ and $\{V^{j}\}$ of a moving system at a specific time steps $t_{i}$ and $t_{j}$ with two LiDAR frames $\{L^{i}_{1}\}$ and $\{L^{j}_{2}\}$, a global reference frame $\{R\}$, and a common calibration frame $\{C\}$.
@@
Fig. 2 shows the spatial relationships, relevant for the calibration task, for a moving vehicle $V$ with two non-overlapping LiDARs $L_{1}$ and $L_{2}$. In addition, we introduce the time step $t_{i}$ for a measurement of $L_{1}$ and the time step $t_{j}$ for a measurement of $L_{2}$ with $1\leq i,j\leq N$. $\{L_{1}^{i}\}$, $\{L_{2}^{j}\}$, $\{V^{i}\}$, and $\{V^{j}\}$ describe the respective LiDAR and vehicle frames at the time steps $t_{i}$ and $t_{j}$. $\{R\}$ represents the global reference frame and $\{C\}$ the common calibration frame as introduced in [10].
${}_{L_{1}}^{V}T$ and ${}_{L_{2}}^{V}T$ describe the ground truth transformations of the LiDARs relative to $\{V\}$. Correspondingly, ${}_{L_{1}}^{V_{1}}T$ and ${}_{L_{2}}^{V_{2}}T$ are the initial transformations that result in the erroneous vehicle frames $\{V_{1}^{i}\}$ and $\{V_{2}^{j}\}$ when applied to the sensors’ point clouds. The S2V calibration errors of $L_{1}$ and $L_{2}$ are described by ${}_{V}^{V_{1}}T$ and ${}_{V}^{V_{2}}T$ respectively.
${}_{V^{i}}^{R}T$ and ${}_{V^{j}}^{R}T$ represent the vehicle poses at time steps $t_{i}$ and $t_{j}$. Applying these transformations to the erroneous vehicle frames $\{V_{1}^{i}\}$ and $\{V_{2}^{j}\}$ results in the erroneous versions $\{R_{1}^{i}\}$ and $\{R_{2}^{j}\}$ of the reference frame, that describe the propagation of a S2V error through the vehicle’s motion. A registration algorithm is executed on point clouds in the frames $\{R_{1}^{i}\}$ and $\{R_{2}^{j}\}$, which returns the calibration transformations ${}^{C}_{R_{1}^{i}}T$ and ${}^{C}_{R_{2}^{j}}T$ relative to the calibration frame. ${}^{C}_{R}T$ represents the calibration frame error relative to the reference frame $\{R\}$. In this specialized dual-LiDAR version of the calibration problem, ${}^{C}_{R}T$ can be computed as follows:
$\begin{split}{}_{R}^{C}T={}_{R_{a}^{i}}^{C}T~{}{}_{V^{i}}^{R}T~{}{}^{V_{a}}_{V}T~{}{}_{V^{i}}^{R}T^{-1}\quad\\ \forall i,a\quad s.t.\quad 1\leq i\leq N,~{}a\in\{1,2\}.\end{split}$ (1)
Further, we define $L_{2}$ to be the reference LiDAR and ${}^{V_{2}}_{V}T$ to be the joint S2V error. The S2S calibration error of $L_{1}$ relative to $L_{2}$ is referred to as ${}_{V_{1}}^{V_{2}}T$. The $L_{1}$ S2V error is therefore defined as follows:
${}^{V_{1}}_{V}T={}_{V_{1}}^{V_{2}}T^{-1}~{}{}_{V}^{V_{2}}T.$ (2)
@@
Since the S2V errors ${}^{V_{a}}_{V}T$ are not known, the ground truth value of ${}^{C}_{R}T$ cannot be computed directly. Therefore, we introduce ${}^{C}_{R}\bar{T}_{a}^{i}$ which describes the approximation of ${}^{C}_{R}T$ for the LiDAR $L_{a}$ at time step $t_{i}$ and an estimated S2V error ${}^{V_{a}}_{V}\bar{T}$. In a non-overlapping setup, the respective LiDARs may not see the target at all time steps. Therefore, we define a set $S$ as the set of all pairs of sensors and time steps for which such a transformation is given, i. e. the sensor sees the target. Specifically, we say $(i,a)\in S$ iff ${}_{R_{a}^{i}}^{C}T$ is known. The problem of recovering the extrinsic sensor parameters can now be reduced to the following minimization problem:
$({}_{V}^{V_{2}}T,~{}{}_{V_{1}}^{V_{2}}T)\approx\mathop{\arg\min}\limits_{({}_{V}^{V_{2}}\bar{T},~{}{}_{V_{1}}^{V_{2}}\bar{T})}\sum\limits_{\{(i,a),~{}(j,b)\}\subset S}e({}_{R}^{C}\bar{T}^{i}_{a},~{}{}_{R}^{C}\bar{T}^{j}_{b}),$ (3)
where $e$ is a function returning the error between a pair of transformations. ${}_{V}^{V_{2}}\bar{T}$ and ${}{}_{V_{1}}^{V_{2}}\bar{T}$ correspond to the S2S and S2V errors and are optimized by the minimization algorithm. The optimal transformations ${}_{V}^{V_{2}}T$ and ${}_{V_{1}}^{V_{2}}T$ are finally used to recover the sensor extrinsics ${}^{V}_{L_{1}}T$ and ${}^{V}_{L_{2}}T$:
${}^{V}_{L_{1}}T={}^{V_{1}}_{V}T^{-1}~{}{}^{V_{1}}_{L_{1}}T={}^{V_{2}}_{V}T^{-1}~{}{}^{V_{2}}_{V_{1}}T~{}{}^{V_{1}}_{L_{1}}T$ (4)
${}^{V}_{L_{2}}T={}^{V_{2}}_{V}T^{-1}~{}{}^{V_{2}}_{L_{2}}T,$ (5)
with ${}^{V_{1}}_{L_{1}}T$ and ${}^{V_{2}}_{L_{2}}T$ being the initial transformations of both LiDAR sensors given as an input to the algorithm.
III-B Implementation Details
@@
We now present the implementation details of our framework used to solve the minimization problem formulated in Section III-A. Specifically, we discuss the kinematic model used to recover the vehicle poses ${}_{V^{i}}^{R}T$, the joint registration algorithm used to obtain the transformations ${}_{R_{a}^{i}}^{C}T$ of all point clouds to a common calibration frame $\{C\}$, as well as the minimization algorithm.
III-B1 Kinematic Model
@@
The approach is compatible with any kinematic model, provided that the vehicle poses are given at the time steps of the LiDAR measurements. Two sources of vehicle poses were used to validate the performance of the algorithm: The ground truth vehicle poses obtained from simulation and an unscented Kalman filter (UKF) [27]. The latter kinematic model is used to combine IMU, GPS, velocity and orientation data to recover vehicle pose estimates ${}_{V^{i}}^{R}T$ at all time steps $t_{i}$. The UKF also allows the modeling of sensor-specific noise and uncertainties to evaluate the calibration accuracy under noisy sensor data.
III-B2 Joint Registration
@@
Before a point cloud registration algorithm can be applied to the point clouds, they must be pre-processed. First, the point clouds are transformed into the global frames $\{R_{1}^{i}\}$ and $\{R_{2}^{j}\}$ with the vehicle pose estimates belonging to the respective measurement timestamps. Then, we separate the target points from the ground points by using the Random Sample Consensus (RANSAC) [28] algorithm to identify and filter out the ground plane. Finally, we use GMMCalib [10] for joint registration of the filtered point clouds into a common calibration frame $\{C\}$. GMMCalib returns the registration transform ${}_{R_{1}^{i}}^{C}T$ and ${}_{R_{2}^{i}}^{C}T$ for each point cloud needed for the optimization procedure. Furthermore, it provides the reconstructed target shape in the calibration frame in the form of a point cloud.
III-B3 Minimization Algorithm
@@
To implement a minimization algorithm, it is necessary to define a cost function to be minimized. As shown in Eq. 3, the cost function is formulated as the sum of all pairwise errors between the estimated calibration frame errors. Note that the previously discussed problem statement assumes that the vehicle poses and calibration transformations are accurate. Therefore, errors introduced by the registration algorithm or the kinematic model may negatively affect the calibration result. To account for such erroneous approximations, we introduce an outlier filter that removes the most extreme outliers of the computed errors. For this purpose, we introduce the parameter $k$. After computing all pairwise errors between calibration frame estimates, we disregard all errors that are below the $k^{th}$ percentile or above the $(100-k)^{th}$ percentile. We set $k$ to 10 in our implementation, as this proved to produce the best results for both ground truth and UKF poses.
@@
The error function $e$ introduced in Eq. 3 computes the error between two estimates ${}_{R}^{C}\bar{T}^{i}_{a}$ and ${}_{R}^{C}\bar{T}^{j}_{b}$ of the calibration frame error transformation. In our implementation, we distinguish between the rotational and translational parts of the inverse of both transformations. Inverting the transformations allows to apply them to a set of points in the calibration frame $\{C\}$. To calculate the error between two transformations, their rotational parts are applied to such a set of points. The errors of the rotational parts are the squared pairwise distances between corresponding transformed points. The additional translation error is simply the squared pairwise difference. As GMMCalib returns the reconstructed shape of the target object, we define the points of the oriented bounding box of this shape to be the point set for this error function.

### v2

@@ Abstract
In autonomous systems, sensor calibration is essential for safe and efficient navigation in dynamic environments. Accurate calibration is a prerequisite for reliable perception and planning tasks such as object detection and obstacle avoidance. Many existing LiDAR calibration methods require overlapping fields of view, while others use external sensing devices or postulate a feature-rich environment. In addition, Sensor-to-Vehicle calibration is not supported by the vast majority of calibration algorithms. In this work, we propose a novel target-based technique for extrinsic Sensor-to-Sensor and Sensor-to-Vehicle calibration of multi-LiDAR systems called CaLiV. This algorithm works for non-overlapping fields of view and does not require any external sensing devices. First, we apply motion to produce field of view overlaps and utilize a simple Unscented Kalman Filter to obtain vehicle poses. Then, we use the Gaussian mixture model-based registration framework GMMCalib to align the point clouds in a common calibration frame. Finally, we reduce the task of recovering the sensor extrinsics to a minimization problem. We show that both translational and rotational Sensor-to-Sensor errors can be solved accurately by our method. In addition, all Sensor-to-Vehicle rotation angles can also be calibrated with high accuracy. We validate the simulation results in real-world experiments. The code is open-source and available on https://github.com/TUMFTM/CaLiV.
@@ I INTRODUCTION
A similar problem occurs in sensor calibration when dealing with non-overlapping sensor setups. The calibration becomes significantly more challenging with sensors that do not share a common field of view (FoV) and, therefore, never simultaneously observe the same features. This is common in applications involving large vehicles, such as buses or trains. Accordingly, non-overlapping sensor calibration frameworks also use motion-based methods, including SLAM or hand-eye calibration [2].
SLAM-based calibration methods generate a map of the environment through point cloud matching and implicitly solve for the extrinsic parameters. These approaches rely heavily on a feature-rich environment, which may not always be available. Hand-eye calibration methods are less dependent on specific environmental features, but do not achieve the calibration accuracy of SLAM algorithms [3]. In addition, these motion-based methods are commonly presented as targetless calibration techniques [4, 5]. Although these targetless approaches offer greater flexibility, they still do not match the accuracy of target-based solutions [6]. In contrast, target-based S2V or non-overlapping S2S calibration has mainly been addressed using external sensing devices [7, 8, 9].
Figure 1: CaLiV: Data is collected by driving a curved trajectory. The data is first aligned using vehicle poses and different perspectives of the calibration target. The S2S calibration is first computed by the algorithm in the optimization phase, and it subsequently serves as the basis for the S2V optimization. The input to the S2S and S2V optimization is a rough point cloud alignment and the shape of the reconstructed target, both generated using GMMCalib [10].
In this work, we present CaLiV, a target-based approach to address the problems of S2V calibration for an arbitrary sensor setup without any additional external sensing devices. CaLiV first performs (non-overlapping) S2S calibration and then utilizes the S2S calibration to compute the S2V calibration parameters. As the underlying alignment model, we rely on GMMCalib [10], which provides a specific calibration frame formulation that we utilize for simultaneous optimization of the S2V calibration parameters. The approach is illustrated in Fig. 1. Our contributions are as follows:
  • To the best of our knowledge, we are the first to present a target-based Sensor-to-Vehicle calibration framework for a multi-LiDAR sensor setup without any external sensing devices.
  • In addition, our approach can be used for non-overlapping Sensor-to-Sensor calibration.
  • We use the GMMCalib formulation with a common calibration frame and generalize it to enable Sensor-to-Vehicle calibration.
@@ II-B Target-Based Calibration
The state-of-the-art shows that for the above requirements, target-based approaches outperform targetless approaches in terms of calibration accuracy and they are not dependent on a feature-rich environment. However, this is currently only achieved with the use of external sensing devices and very specific calibration targets, which in some cases, as shown in [8], are even measured in advance. Accuracy is therefore achieved at the expense of the flexibility that can be provided by targetless approaches. We conclude that there is no target-based, environment-independent calibration framework for extrinsic S2V calibration of multi-LiDAR systems with non-overlapping FoVs. Therefore, we propose a target-based calibration framework that enables both S2V and non-overlapping calibration. We achieve this without the use of additional external sensor systems.
@@ III METHODOLOGY
This section describes the methodology behind the framework of Fig. 1. For the sake of clarity, we consider the calibration with a vehicle with two non-overlapping LiDARs. However, the proposed technique can easily be generalized for any multi-LiDAR setup.
III-A Spatial Relationships
@@
Figure 2: Spatial relationship between the frames $\{V^{i}\}$ of a moving system at a specific time steps $t_{i}$ with two LiDAR frames $\{L^{i}_{1}\}$ and $\{L^{i}_{2}\}$, a global reference frame $\{R\}$, and a common calibration frame $\{C\}$.
@@
Fig. 2 shows the spatial relationships, relevant for the calibration task, for a moving vehicle $V$ with two non-overlapping LiDARs $L_{1}$ and $L_{2}$. $\{L_{1}^{i}\}$, $\{L_{2}^{i}\}$ and $\{V^{i}\}$ describe the respective LiDAR and vehicle frames at a time step $t_{i}$ with $1\leq i\leq N$. $\{R\}$ represents the global reference frame and $\{C\}$ the common calibration frame as introduced in [10].
$_{L_{1}}^{V}T$ and $_{L_{2}}^{V}T$ describe the ground-truth transformations of the LiDARs relative to $\{V^{i}\}$ at all time steps. Recovering these ground-truth values is the goal in our calibration. Initially, we are only given erroneous LiDAR transformations, which will be referred to as $_{L_{1}}^{V}\hat{T}$ and $_{L_{2}}^{V}\hat{T}$ in the following. Further, $_{V^{i}}^{R}T$ represents the vehicle pose at time step $t_{i}$ relative to the global reference frame and must be known before the calibration.
In addition to the relationships depicted in Fig. 2, we choose $L_{2}$ to be the reference sensor in our setup. Therefore, we define the transformation $^{L_{2}}_{L_{1}}T$ to be the ground truth S2S transformation and $^{L_{2}}_{L_{1}}\hat{T}$ the initial S2S transformation. The S2V transformation of $L_{1}$ can then be defined as:
$^{V}_{L_{1}}T=^{V}_{L_{2}}T~^{L_{2}}_{L_{1}}T$ (1)
III-B Rough Point Cloud Alignment
As introduced in GMMCalib [10], the point clouds are not aligned by matching source and target point clouds but rather by registering point clouds of different frames to a latent calibration frame. Accordingly, $_{L^{i}_{1}}^{C}T$ and $_{L^{i}_{2}}^{C}T$ are transformations that align the point clouds of a LiDAR at a specific time step in the common calibration frame $\{C\}$.
Before applying GMMCalib, we perform a pre-alignment step by transforming the point clouds with the initial LiDAR transformations $_{L_{1}}^{V}\hat{T}$ and $_{L_{2}}^{V}\hat{T}$, respectively, as well as the ego vehicle pose $_{V^{i}}^{R}T$ at the measurement’s time step. In total, the alignment transformations for the LiDARs are defined by the following equation:
$\begin{split}{}_{L^{i}_{a}}^{C}T={}_{L^{i}_{a}}^{C}T^{{}^{\prime}}~{}_{V^{i}}^{R}T~{}_{L_{a}}^{V}\hat{T}\quad\quad\quad\\ \forall i,a\quad s.t.\quad 1\leq i\leq N,~a\in\{1,2\}.\end{split}$ (2)
$_{L^{i}_{a}}^{C}T^{{}^{\prime}}$ here is the part of the alignment that is done by the point cloud registration algorithm of GMMCalib and $a$ represents the specific LiDAR sensor.
III-C Non-Linear Optimization
$^{C}_{R}T$ represents the pose of the reference frame $\{R\}$ relative to the calibration frame $\{C\}$. Note that this latent frame formulation allows us to compute the Sensor-to-Vehicle calibration in the first place by assuming a fixed transformation between the calibration frame and the reference frame. In this specialized dual-LiDAR version of the calibration problem, $^{C}_{R}T$ can be computed as follows:
$\begin{split}{}_{R}^{C}T={}_{L_{a}^{i}}^{C}T~{}_{L_{a}}^{V}T^{-1}~{}_{V^{i}}^{R}T^{-1}\quad\quad\\ \forall i,a\quad s.t.\quad 1\leq i\leq N,~a\in\{1,2\}.\end{split}$ (3)
This simply follows by the relationships depicted in Fig. 2.
Since the LiDARs’ ground truth S2V transformation $_{L_{2}}^{V}T$ and S2S transformation $_{L_{1}}^{L_{2}}T$ are unknown, we cannot compute $_{R}^{C}T$ directly. Therefore, we introduce $^{C}_{R}\hat{T}_{a}^{i}$ which describes the approximation of $^{C}_{R}T$ for the LiDAR $L_{a}$ at time step $t_{i}$ for specific estimates $_{L_{2}}^{V}T^{{}^{\prime}}$ and $_{L_{1}}^{L_{2}}T^{{}^{\prime}}$ of the LiDAR transformations. Note that these estimates can differ from the initial transformations and are chosen in an iterative minimization process. Then, we can define an error function $e({}_{R}^{C}\hat{T}^{i}_{a},{}_{R}^{C}\hat{T}^{j}_{b})$ which returns the error between a pair of approximated transformations between reference and calibration frame.
@@ III-C Non-Linear Optimization
In a non-overlapping setup, the respective LiDARs may not see the target at all time steps and thus we cannot simply sum the results of the error function for all LiDARs at all time steps to evaluate our current LiDAR transformation estimates. Therefore, we define a set $S$ as the set of all pairs of sensors and time steps for which such a transformation is given, i. e. the sensor sees the target. Specifically, we say $(i,a)\in S$ iff $_{L^{i}_{a}}^{C}T$ is known.
Recovering the optimal transformations $_{L_{2}}^{V}T^{\star}$ and $_{L_{1}}^{L_{2}}T^{\star}$ can now be reduced to the following minimization problem:
$_{L_{2}}^{V}T^{\star},{}_{L_{1}}^{L_{2}}T^{\star}=\operatorname*{arg\,min}_{{}_{L_{2}}^{V}T^{{}^{\prime}},{}_{L_{1}}^{L_{2}}T^{{}^{\prime}}}\ \sum_{\{(i,a),~(j,b)\}\subset S}e\!\left({}_{R}^{C}\hat{T}^{i}_{a},~{}_{R}^{C}\hat{T}^{j}_{b}\right).$ (4)
This same minimization problem can then be formulated with a fixed S2S transformation $_{L_{1}}^{L_{2}}T^{\star}$ in a second iteration, which further improved S2V calibration results in our experiments.
III-D Implementation Details
@@
We now present the implementation details of our framework used to solve the minimization problem formulated in Section III-C. Specifically, we discuss the kinematic model used to recover the vehicle poses $_{V^{i}}^{R}T$, the joint registration algorithm used to obtain the transformations $_{L^{i}_{a}}^{C}T^{{}^{\prime}}$ of all point clouds to a common calibration frame $\{C\}$, as well as the minimization algorithm.
III-D1 Kinematic Model
@@
The approach is compatible with any kinematic model, provided that the vehicle poses are given at the time steps of the LiDAR measurements. We use an Unscented Kalman Filter (UKF) for the kinematic model that combines IMU, GPS, velocity and orientation data to recover vehicle pose estimates $_{V^{i}}^{R}T$ at all time steps $t_{i}$. Our UKF also allows the modeling of sensor-specific noise and uncertainties to evaluate the calibration accuracy under noisy sensor data.
III-D2 Joint Registration
@@
Before a point cloud registration algorithm can be applied to the point clouds, they must be pre-processed. First, the point clouds are transformed with the initial LiDAR transformations and the ego vehicle pose at the measurement’s time step, as discussed in Section III-B. Then, we separate the target points from the ground points by using the Random Sample Consensus (RANSAC) [26] algorithm to identify and filter out the ground plane. Finally, we use GMMCalib [10] for joint registration of the filtered point clouds into a common calibration frame $\{C\}$. GMMCalib returns the registration transform $_{L^{i}_{1}}^{C}T^{{}^{\prime}}$ and $_{L^{i}_{2}}^{C}T^{{}^{\prime}}$ for each point cloud that contains points belonging to the target shape.
III-D3 Minimization Algorithm
@@
As shown in Eq. 4, the cost function is formulated as the sum of all pairwise errors of the estimated transformations between calibration frame and reference frame. Note that the previously discussed problem statement assumes that the vehicle poses and calibration transformations are accurate. To account for errors introduced by the registration algorithm or the kinematic model, we introduce an outlier filter that removes the most extreme outliers of the computed errors. For this purpose, we introduce the parameter $k$. After computing all pairwise errors between estimates of $_{R}^{C}T$, we disregard all errors that are below the $k^{th}$ percentile or above the $(100-k)^{th}$ percentile. We set $k$ to 10 in our implementation, as this proved to produce the best results for our experiments.
@@
Figure 3: This figure illustrates the optimization process of the calibration. The two transformation estimates are used to transform the point sets represented by the bounding box of the calibration target shape.
The error function $e$ introduced in Section III-C computes the error between two estimates $_{R}^{C}\hat{T}^{i}_{a}$ and $_{R}^{C}\hat{T}^{j}_{b}$ of the transformation between calibration frame and reference frame. We invert the transformations to be able to apply them to a set of points in the calibration frame $\{C\}$. To calculate the error between two transformations, we transform the set of points and compute the their squared pairwise distances. With GMMCalib returning the reconstructed shape of the target object, we define the points of the oriented bounding box of this shape to be the point set for this error function as illustrated in Figure 3. In contrast to the complete reconstructed target point cloud, the bounding box representation of the point set decreases the performance overhead in non-linear optimization.